OpenStudy (anonymous):

5x+ 20=7x need help with this question

OpenStudy (saifoo.khan):

x = 10

OpenStudy (anonymous):

How

OpenStudy (saifoo.khan):

7x - 5x = 20 2x = 20 x = 20/2 x = 10

OpenStudy (anonymous):

Thanks

OpenStudy (saifoo.khan):

Welks/

OpenStudy (saifoo.khan):

Please Hit good answer.